Cu(s) + AgNO3(aq) ---> CuNO3(aq) + Ag(s)

How many grams of silver would be produced if you began with 10 grams of copper?


1
Expert's answer
2020-12-24T03:53:58-0500

MM(Cu) = 63.5 g/mol

"n(Cu) = \frac{10}{63.5} = 0.157 \;mol"

n(Ag) = n(Cu) = 0.157 mol

MM(Ag) = 107.8 g/mol

"m(Ag) = 0.157 \times 107.8 = 16.92 \;g"

Answer: 16.92 g